Adventure Travel signed a 9​%, 10−year note for $164,000. The company paid an installment of $2,500

for the first month. After the first​ payment, what is the principal​ balance?

A. ) $ 166500

B. ) $162730

C. ) $161500

D. ) 1627700
Answer #1

Answer : B) $162,730

interest for the First Month = 164,000*9%*1/12 = $1,230

Principal amount in the Installment = 2,500-1,230= $1,270

Principal Balance = 164,000-1,270 = $162,730(Answer)